# Cosmos World Foundation Model Platform for Physical AI

### v1

@@ Abstract
Physical AI needs to be trained digitally first. It needs a digital twin of itself, the policy model, and a digital twin of the world, the world model. In this paper, we present the Cosmos World Foundation Model Platform to help developers build customized world models for their Physical AI setups. We position a world foundation model as a general-purpose world model that can be fine-tuned into customized world models for downstream applications. Our platform covers a video curation pipeline, pre-trained world foundation models, examples of post-training of pre-trained world foundation models, and video tokenizers. To help Physical AI builders solve the most critical problems of our society, we make our platform open-source and our models open-weight with permissive licenses available via NVIDIA Cosmos.
@@ 1 Introduction
In this paper, we introduce the Cosmos World Foundation Model (WFM) Platform for building Physical AI. We are mainly concerned with the visual world foundation model, where the observations are presented as videos, and the perturbations can exist in various forms. As illustrated in Fig. 2, we present a pre-training-and-then-post-training paradigm, where we divide WFMs into pre-trained and post-trained WFMs. To build a pre-trained WFM, we leverage a large-scale video training dataset to expose the model to a diverse set of visual experiences so it can become a generalist. To build a post-trained WFM, we fine-tune the pre-trained WFM to arrive at a specialized WFM using a dataset collected from a particular Physical AI environment for the targeted, specialized Physical AI setup. Fig. 1 shows example results from our pre-trained and post-trained WFMs.
@@
We aim to build a world foundation model platform to help Physical AI builders advance their systems. To achieve this goal, we make our pre-trained world foundation models and tokenizers available under the NVIDIA Open Model License at NVIDIA Cosmos and NVIDIA Cosmos Tokenizer respectively. The pre-training script and post-training script will be available at NVIDIA Nemo Framework together with the video data curation pipeline to help builders craft their fine-tuning datasets. While this paper makes several improvements in world foundation model design, the world foundation model problem is still far from being solved. Additional research is required to advance the state-of-the-art further.
@@ 2.2 Current Cosmos
Fig. 4 visualizes what is available in the Cosmos WFM platform that is included in this paper, which includes video curator, video tokenization, world foundation model pre-training, world foundation model post-training, and guardrail.
@@ 4 Tokenizer
As shown in Fig. 8, our evaluation results demonstrate Cosmos Tokenizer significantly outperforms existing tokenizers by a large margin—for instance, achieving a +4 dB PSNR improvement in reconstruction quality on DAVIS videos. It runs up to faster and can encode videos up to 8 seconds at 1080p and 10 seconds at 720p in one shot without running out of memory on a single NVIDIA A100 GPU with 80GB memory. A suite of pre-trained models, with spatial compression of and , and temporal compression factors of and is available at github.com/NVIDIA/Cosmos-Tokenizer.
@@ 4.2 Training Strategy
Cosmos-1.0-Tokenizer: Trained using mini-batches sampling a larger number of video frames (121 frames for CV and 49 frames for DV).
@@ 5 World Foundation Model Pre-training
| Type | Models | Tokenizer | Enhancer |
| Diffusion | Cosmos-1.0- Diffusion-7B- Text2World Cosmos-1.0- Diffusion-7B- Video2World | Cosmos-1.0- Tokenizer- CV8x8x8 | Cosmos-1.0- PromptUpsampler- 12B-Text2World |
| Cosmos-1.0- Diffusion-14B- Text2World Cosmos-1.0- Diffusion-14B- Video2World | |||
| Autoregressive | Cosmos-1.0- Autoregressive- 4B Cosmos-1.0- Autoregressive- 5B-Video2World | Cosmos-1.0- Tokenizer- DV8x16x16 | Cosmos-1.0- Diffusion-7B- Decoder- DV8x16x16ToCV8x8x8 |
| Cosmos-1.0- Autoregressive- 12B Cosmos-1.0- Autoregressive- 13B-Video2World |
In Tab. 10, we present a map of our pre-trained WFMs and their companions. For the diffusion-based WFM family, we start by building two Text2World models of 7B and 14B, respectively, which render Cosmos-1.0-Diffusion-7B-Text2World and Cosmos-1.0-Diffusion-14B-Text2World. These models can map text prompts to videos of visual worlds. We then fine-tune the Text2World models to take additional video input, representing the current observation. The result is a Video2World model where the future video is predicted based on the current observation (input video) and the perturbation (text prompt). These diffusion models are latent diffusion models that take continuous tokens. We use Cosmos-1.0-Tokenizer-CV8x8x8 to produce the visual tokens. The training text prompts for the WFMs are produced by a VLM through video description generation. These descriptions follow a different distribution of human descriptions of videos. To mitigate the domain gap, we build Cosmos-1.0-PromptUpsampler-12B-Text2World based on the Mistral-NeMo-12B-Instruct model (Mistral and NVIDIA, 2024) to help convert human text prompts to those preferred by our diffusion-based WFMs.
For the autoregressive-based WFM family, we first build two base models that are 4B and 12B in size, respectively, to predict future videos purely based on the current video observation. We name them Cosmos-1.0-Autoregressive-4B and Cosmos-1.0-Autoregressive-12B, respectively. These are Llama3-style GPT models trained from scratch for the video prediction task and bear no language understanding. To enable autoregressive-based WFMs to utilize textual information for next token prediction, we incorporate T5 embeddings of the input text prompt into the WFMs through cross-attention layers added to the transformer blocks. These autoregressive WFMs use Cosmos-1.0-Tokenizer-DV8x16x16, which maps an input video to a few integers. The heavy compression of the tokenizer can sometimes lead to undesired distortions. To address the problem, we build a diffusion decoder (Cosmos-1.0-Diffusion-7B-Decoder-DV8x16x16ToCV8x8x8) through fine-tuning the Cosmos-1.0-Diffusion-7B-Text2World model to map discrete tokens in the DV8x16x16 space to continuous tokens in the CV8x8x8 space.
@@ 5.1 Diffusion-based World Foundation Model
Our diffusion-based WFMs are latent diffusion models that operate within a learned latent space of a tokenizer, enabling a compact, reduced-dimensional representation of videos. This design choice offers several advantages: it reduces computational costs during both training and inference while simplifying the denoising task (Rombach et al., 2022; Hoogeboom et al., 2024). To tokenize videos into latent representations, we employ Cosmos-1.0-Tokenizer-CV8x8x8.
@@ 5.1.2 Architecture
AdaLN-LoRA. We find that DiT’s adaptive layer normalization (AdaLN) layers (Xu et al., 2019; Peebles and Xie, 2023) account for a significant portion of the model parameters while contributing negligibly to the computational complexity in terms of FLOPs. Inspired by W.A.L.T (Gupta et al., 2024a), we implement Low-Rank Adaptation (LoRA) (Hu et al., 2022) to decompose the dense linear projections in these layers into low-rank approximations. For Cosmos-1.0-Diffusion-7B, this architectural optimization achieves a 36% reduction in parameter count (from 11B to 7B parameters) while maintaining performance parity across all evaluation metrics, demonstrating the effectiveness of our parameter-efficient design.
@@ 5.1.4 Scaling Up
For instance, our 14B model (Cosmos-1.0-Diffusion-14B-Text2World) requires approximately 280 GB for model parameters, gradients, and optimizer states, alongside 310 GB for activations during high-resolution pre-training. Given the 80GB HBM3 limit of NVIDIA H100 GPUs, we employ Fully Sharded Data Parallelism (FSDP) and Context Parallelism (CP) to distribute memory demands across multiple GPUs.
@@
Using Cosmos-1.0-Diffusion-14B as an instance, employing FSDP with a sharding factor of 64 reduces memory requirements for parameters, gradients, and optimizer states, bringing them down from 280 GB to approximately . Similarly, employing CP with decreases activation memory from 310 GB to roughly . It is important to note that these calculations are underestimations; in practice, additional memory is consumed by the tokenizer and unsharded parameters. Overlapping communication and computation in CP also necessitates each GPU to retain multiple chunks of .
@@ 5.1.5 Prompt Upsampler
Prompt upsampler for Text2World model. We fine-tune Mistral-NeMo-12B-Instruct (Mistral and NVIDIA, 2024) to build our prompt upsampler. To obtain paired data, that is, short prompts simulating user input and the corresponding long prompts reflecting the distribution of training prompts, we use a VLM to generate short captions based on our training long prompts and corresponding videos. This long-to-short data creation strategy is effective in (1) preserving the authentic video content and distribution from detailed training prompts of WFMs and (2) ensuring fidelity between the short and long prompts. The resulting prompt upsampler is termed Cosmos-1.0-PromptUpsampler-12B-Text2World.
@@ 5.1.6 Results
In Fig. 12, we present qualitative results generated by our Cosmos-1.0-Diffusion-7B-Text2World and Cosmos-1.0-Diffusion-14B-Text2World models. Both models produce videos of high visual quality, motion dynamics, and text alignment. Compared to the 7B model, the 14B model is able to generate videos capturing more complex visual details and intricate motions.
@@ 5.2.1 Architecture
For our autoregressive models, we use our Cosmos-1.0-Tokenizer-DV8x16x16 as the tokenizer. As introduced in Sec. 4, we leverage the Finite-Scalar-Quantization (FSQ) (Mentzer et al., 2023) to quantize the -dimensional latent space into levels. This quantization leads to a vocabulary size of .
@@ 5.2.2 Scaling Up
For instance, our 12B model (Cosmos-1.0-Autoregressive-12B) demands approximately 192 GB of memory for its parameters, gradients, and optimizer states combined. As this is beyond a single NVIDIA H100 GPU’s 80GB HBM3 capacity, we leverage tensor parallelism (TP) (Shoeybi et al., 2019) and its extension, sequence parallelism (SP) (Korthikanti et al., 2023), to distribute the memory requirements and computation across multiple GPUs.
@@ 5.2.3 Training Strategy
Cosmos-1.0-Autoregressive-4B: a 4B transformer model for next video token prediction. This model is trained using stage 1 and stage 1.1 of the multi-stage training objective.
Cosmos-1.0-Autoregressive-5B-Video2World: a 5B transformer model derived from our Cosmos-1.0-
Autoregressive-4B and trained additionally with stage 2 of the multi-stage training objective. -
Cosmos-1.0-Autoregressive-12B: a 12B transformer model for next video token prediction. This model is trained using stage 1 and stage 1.1 of the multi-stage training objective.
Cosmos-1.0-Autoregressive-13B-Video2World: a 13B transformer model derived from Cosmos-1.0-Autoregressive-12B and trained additionally with stage 2 of the multi-stage training objective.
@@ 5.2.4 Inference Optimization Towards Real-Time Generation
Low-resolution adaptation for real-time inference. We pursue real-time inference by adapting our model to a lower spatial resolution of , which results in a lower number of tokens per video. Specifically, we first fine-tune the discrete video tokenizer (Cosmos-1.0-Tokenizer-DV81616 in Sec. 4) on 320p low-resolution videos using videos from the target Physical AI domain. Then, we fine-tune our autoregressive WFM that is pre-trained in resolution (Cosmos-1.0-Autoregressive-4B in Sec. 5.2.3) with this low-resolution tokenizer on videos of resolution from the target Physical AI domain. Finally, we add the Medusa heads to the fine-tuned low-resolution autoregressive WFM.
@@
| Cosmos-1.0-Autoregressive-4B (with Medusa) | 806.61 | 10.08 |
@@ 5.2.5 Diffusion Decoder
Our Cosmos tokenizer uses a lightweight encoder-decoder architecture to perform aggressive compression, which reduces the number of tokens for our WFM training. As a result of aggressive compression, it could sometimes lead to blurriness and visible artifacts in video generation, especially in the autoregressive WFM setting, where only a few integers are used to represent a rich video through discrete tokenization. We resort to the diffusion decoder design (Ramesh et al., 2022; OpenAI, 2024a) to address the limitation. Specifically, we build a more powerful tokenizer decoder by fine-tuning Cosmos-1.0-Diffusion-7B-Text2Video in Sec. 5.1.
Fig. 16 illustrates how we train a diffusion decoder for our autoregressive WFMs. For each training video, we use Cosmos-1.0-Tokenizer-CV8x8x8 and Cosmos-1.0-Tokenizer-DV8x16x16 to compute a continuous token video and a corresponding discrete token video, respectively. We note that Cosmos-1.0-Tokenizer-CV8x8x8 can produce higher quality video outputs than Cosmos-1.0-Tokenizer-DV8x16x16 thanks to the more gentle continuous tokenization process and the less aggressive compression scheme ( instead of ).
The discrete token video is treated as the conditional input to the denoiser of the Cosmos-1.0-Diffusion-7B model. To compute the conditional input, we first embed each discrete token of the discrete token video into a 16-dimensional vector based on a learnable vocabulary embedding layer. We then upsample the embedding along the and directions so that the conditional input will be of the same size as the noisy input to the denoiser from the continuous token video. We concatenate the noisy continuous inputs with the conditional inputs along the channel dimension, which becomes the input to the diffusion denoiser. The first layer of the denoiser is channel-dimension expanded to accommodate the new input shape. We fine-tune the updated Cosmos-1.0-Diffusion-7B by removing the added noise. As the discrete token video is not noise-corrupted, the denoiser learns to leverage the residing information in the conditional input for denoising. The result is a higher-quality decoder for the tokenizer that decodes the discrete token by solving a reserve diffusion problem.
Fig. 16 illustrates the inference. The output discrete token video (under discrete compression) from our autoregressive WFM is decoded into a video through two steps. First, we roll out the conditional denoiser to generate a continuous token video (under continuous compression) based on the autoregressive WFM output. Next, the continuous token video is decoded by Cosmos-1.0-Tokenizer-CV8x8x8 to produce the resulting RGB video.
@@
| Output of Cosmos-1.0-Autoregressive-13B-Video2World | |||
| Output of Cosmos-1.0-Autoregressive-13B-Video2World + diffusion decoder | |||

### v2

@@ Abstract
Physical AI needs to be trained digitally first. It needs a digital twin of itself, the policy model, and a digital twin of the world, the world model. In this paper, we present the Cosmos World Foundation Model Platform to help developers build customized world models for their Physical AI setups. We position a world foundation model as a general-purpose world model that can be fine-tuned into customized world models for downstream applications. Our platform covers a video curation pipeline, pre-trained world foundation models, examples of post-training of pre-trained world foundation models, and video tokenizers. To help Physical AI builders solve the most critical problems of our society, we make Cosmos open-source and our models open-weight with permissive licenses available via NVIDIA Cosmos-Predict1.
@@ 1 Introduction
In this paper, we introduce the Cosmos World Foundation Model (WFM) Platform for building Physical AI. We are mainly concerned with the visual world foundation model, where the observations are presented as videos, and the perturbations can exist in various forms. As illustrated in Fig. 2, we present a pre-training-and-then-post-training paradigm, where we divide WFMs into pre-trained and post-trained WFMs. To build a pre-trained WFM, we leverage a large-scale video training dataset to expose the model to a diverse set of visual experiences to become a generalist. To build a post-trained WFM, we fine-tune the pre-trained WFM to arrive at a specialized WFM using a dataset collected from a particular Physical AI environment for the targeted, specialized Physical AI setup. Fig. 1 shows example results from our pre-trained and post-trained WFMs.
@@
We aim to build a world foundation model platform to help Physical AI builders advance their systems. To achieve this goal, we make our pre-trained world foundation models and tokenizers available under the NVIDIA Open Model License at NVIDIA Cosmos. While this paper makes several improvements in world foundation model design, the world foundation model problem is still far from being solved. Additional research is required to advance the state-of-the-art further.
@@ 2.2 Current Cosmos
Fig. 4 visualizes what is available in the Cosmos WFM platform in this paper, including video curator, video tokenization, world foundation model pre-training, world foundation model post-training, and guardrail.
@@ 4 Tokenizer
As shown in Fig. 8, our evaluation results demonstrate Cosmos Tokenizer significantly outperforms existing tokenizers by a large margin—for instance, achieving a +4 dB PSNR improvement in reconstruction quality on DAVIS videos. It runs up to faster and can encode videos up to 8 seconds at 1080p and 10 seconds at 720p in one shot without running out of memory on a single NVIDIA A100 GPU with 80GB memory.
@@ 4.2 Training Strategy
Cosmos-Tokenize1: Trained using mini-batches sampling a larger number of video frames (121 frames for CV and 49 frames for DV).
@@ 5 World Foundation Model Pre-training
| Type | Diffusion | Autoregressive |
| Models | 1. Cosmos-Predict1- 7B-Text2World Cosmos-Predict1- 7B-Video2World | 1. Cosmos-Predict1- 4B Cosmos-Predict1- 5B-Video2World |
| 2. Cosmos-Predict1- 14B-Text2World Cosmos-Predict1- 14B-Video2World | 2. Cosmos-Predict1- 12B Cosmos-Predict1- 13B-Video2World | |
| Tokenizer | Cosmos-Tokenize1-CV888-720p | Cosmos-Tokenize1-DV81616-720p |
| Enhancer | Cosmos-UpsamplePrompt1-12B-Text2World | Cosmos-Predict1-7B-Decoder- DV81616ToCV888-720p |
In Tab. 10, we present a map of our pre-trained WFMs and their companions. For the diffusion-based WFM family, we start by building two Text2World models of 7B and 14B, respectively, which render Cosmos-Predict1-7B-Text2World and Cosmos-Predict1-14B-Text2World. These models can map text prompts to videos of visual worlds. We then fine-tune the Text2World models to take additional video input, representing the current observation. The result is a Video2World model where the future video is predicted based on the current observation (input video) and the perturbation (text prompt). These diffusion models are latent diffusion models that take continuous tokens. We use Cosmos-Tokenize1-CV888-720p to produce the visual tokens. The training text prompts for the WFMs are produced by a VLM through video description generation. These descriptions follow a different distribution of human descriptions of videos. To mitigate the domain gap, we build Cosmos-UpsamplePrompt1-12B-Text2World based on the Mistral-NeMo-12B-Instruct model (Mistral and NVIDIA, 2024) to help convert human text prompts to those preferred by our diffusion-based WFMs.
For the autoregressive-based WFM family, we first build two base models that are 4B and 12B in size, respectively, to predict future videos purely based on the current video observation. We name them Cosmos-Predict1-4B and Cosmos-Predict1-12B, respectively. These are Llama3-style GPT models trained from scratch for the video prediction task and bear no language understanding. To enable autoregressive-based WFMs to utilize textual information for next token prediction, we incorporate T5 embeddings of the input text prompt into the WFMs through cross-attention layers added to the transformer blocks. These autoregressive WFMs use Cosmos-Tokenize1-DV81616-720p, which maps an input video to a few integers. The heavy compression of the tokenizer can sometimes lead to undesired distortions. To address the problem, we build a diffusion decoder (Cosmos-Predict1-7B-Decoder-DV81616ToCV888-720p) through fine-tuning the Cosmos-Predict1-7B-Text2World model to map discrete tokens in the DV81616 space to continuous tokens in the CV888 space.
@@ 5.1 Diffusion-based World Foundation Model
Our diffusion-based WFMs are latent diffusion models that operate within a learned latent space of a tokenizer, enabling a compact, reduced-dimensional representation of videos. This design choice offers several advantages: it reduces computational costs during both training and inference while simplifying the denoising task (Rombach et al., 2022; Hoogeboom et al., 2024). To tokenize videos into latent representations, we employ Cosmos-Tokenize1-CV888-720p.
@@ 5.1.2 Architecture
AdaLN-LoRA. We find that DiT’s adaptive layer normalization (AdaLN) layers (Xu et al., 2019; Peebles and Xie, 2023) account for a significant portion of the model parameters while contributing negligibly to the computational complexity in terms of FLOPs. Inspired by W.A.L.T (Gupta et al., 2024a), we implement Low-Rank Adaptation (LoRA) (Hu et al., 2022) to decompose the dense linear projections in these layers into low-rank approximations. For Cosmos-Predict1-7B, this architectural optimization achieves a 36% reduction in parameter count (from 11B to 7B parameters) while maintaining performance parity across all evaluation metrics, demonstrating the effectiveness of our parameter-efficient design.
@@ 5.1.4 Scaling Up
For instance, our 14B model (Cosmos-Predict1-14B-Text2World) requires approximately 280 GB for model parameters, gradients, and optimizer states, alongside 310 GB for activations during high-resolution pre-training. Given the 80GB HBM3 limit of NVIDIA H100 GPUs, we employ Fully Sharded Data Parallelism (FSDP) and Context Parallelism (CP) to distribute memory demands across multiple GPUs.
@@
Using Cosmos-Predict1-14B as an instance, employing FSDP with a sharding factor of 64 reduces memory requirements for parameters, gradients, and optimizer states, bringing them down from 280 GB to approximately . Similarly, employing CP with decreases activation memory from 310 GB to roughly . It is important to note that these calculations are underestimations; in practice, additional memory is consumed by the tokenizer and unsharded parameters. Overlapping communication and computation in CP also necessitates each GPU to retain multiple chunks of .
@@ 5.1.5 Prompt Upsampler
Prompt upsampler for Text2World model. We fine-tune Mistral-NeMo-12B-Instruct (Mistral and NVIDIA, 2024) to build our prompt upsampler. To obtain paired data, that is, short prompts simulating user input and the corresponding long prompts reflecting the distribution of training prompts, we use a VLM to generate short captions based on our training long prompts and corresponding videos. This long-to-short data creation strategy is effective in (1) preserving the authentic video content and distribution from detailed training prompts of WFMs and (2) ensuring fidelity between the short and long prompts. The resulting prompt upsampler is termed Cosmos-UpsamplePrompt1-12B-Text2World.
@@ 5.1.6 Results
In Fig. 12, we present qualitative results generated by our Cosmos-Predict1-7B-Text2World and Cosmos-Predict1-14B-Text2World models. Both models produce videos of high visual quality, motion dynamics, and text alignment. Compared to the 7B model, the 14B model is able to generate videos capturing more complex visual details and intricate motions.
@@ 5.2.1 Architecture
For our autoregressive models, we use our Cosmos-Tokenize1-DV81616-720p as the tokenizer. As introduced in Sec. 4, we leverage the Finite-Scalar-Quantization (FSQ) (Mentzer et al., 2023) to quantize the -dimensional latent space into levels. This quantization leads to a vocabulary size of .
@@ 5.2.2 Scaling Up
For instance, our 12B model (Cosmos-Predict1-12B) demands approximately 192 GB of memory for its parameters, gradients, and optimizer states combined. As this is beyond a single NVIDIA H100 GPU’s 80GB HBM3 capacity, we leverage tensor parallelism (TP) (Shoeybi et al., 2019) and its extension, sequence parallelism (SP) (Korthikanti et al., 2023), to distribute the memory requirements and computation across multiple GPUs.
@@ 5.2.3 Training Strategy
Cosmos-Predict1-4B: a 4B transformer model for next video token prediction. This model is trained using stage 1 and stage 1.1 of the multi-stage training objective.
Cosmos-Predict1-5B-Video2World: a 5B transformer model derived from our Cosmos-Predict1-4B and trained additionally with stage 2 of the multi-stage training objective.
Cosmos-Predict1-12B: a 12B transformer model for next video token prediction. This model is trained using stage 1 and stage 1.1 of the multi-stage training objective.
Cosmos-Predict1-13B-Video2World: a 13B transformer model derived from Cosmos-Predict1-12B and trained additionally with stage 2 of the multi-stage training objective.
@@ 5.2.4 Inference Optimization Towards Real-Time Generation
Low-resolution adaptation for real-time inference. We pursue real-time inference by adapting our model to a lower spatial resolution of , which results in a lower number of tokens per video. Specifically, we first fine-tune the discrete video tokenizer (Cosmos-Tokenize1-DV81616-720p in Sec. 4) on 320p low-resolution videos using videos from the target Physical AI domain. Then, we fine-tune our autoregressive WFM that is pre-trained in resolution (Cosmos-Predict1-4B in Sec. 5.2.3) with this low-resolution tokenizer on videos of resolution from the target Physical AI domain. Finally, we add the Medusa heads to the fine-tuned low-resolution autoregressive WFM.
@@
| Cosmos-Predict1-4B (with Medusa) | 806.61 | 10.08 |
@@ 5.2.5 Diffusion Decoder
Our Cosmos tokenizer uses a lightweight encoder-decoder architecture to perform aggressive compression, which reduces the number of tokens for our WFM training. As a result of aggressive compression, it could sometimes lead to blurriness and visible artifacts in video generation, especially in the autoregressive WFM setting, where only a few integers are used to represent a rich video through discrete tokenization. We resort to the diffusion decoder design (Ramesh et al., 2022; OpenAI, 2024a) to address the limitation. Specifically, we build a more powerful tokenizer decoder by fine-tuning Cosmos-Predict1-7B-Text2Video in Sec. 5.1.
Fig. 16 illustrates how we train a diffusion decoder for our autoregressive WFMs. For each training video, we use Cosmos-Tokenize1-CV888-720p and Cosmos-Tokenize1-DV81616-720p to compute a continuous token video and a corresponding discrete token video, respectively. We note that Cosmos-Tokenize1-CV888-720p can produce higher quality video outputs than Cosmos-Tokenize1-DV81616-720p thanks to the more gentle continuous tokenization process and the less aggressive compression scheme ( instead of ).
The discrete token video is treated as the conditional input to the denoiser of the Cosmos-Predict1-7B model. To compute the conditional input, we first embed each discrete token of the discrete token video into a 16-dimensional vector based on a learnable vocabulary embedding layer. We then upsample the embedding along the and directions so that the conditional input will be of the same size as the noisy input to the denoiser from the continuous token video. We concatenate the noisy continuous inputs with the conditional inputs along the channel dimension, which becomes the input to the diffusion denoiser. The first layer of the denoiser is channel-dimension expanded to accommodate the new input shape. We fine-tune the updated Cosmos-Predict1-7B by removing the added noise. As the discrete token video is not noise-corrupted, the denoiser learns to leverage the residing information in the conditional input for denoising. The result is a higher-quality decoder for the tokenizer that decodes the discrete token by solving a reserve diffusion problem.
Fig. 16 illustrates the inference. The output discrete token video (under discrete compression) from our autoregressive WFM is decoded into a video through two steps. First, we roll out the conditional denoiser to generate a continuous token video (under continuous compression) based on the autoregressive WFM output. Next, the continuous token video is decoded by Cosmos-Tokenize1-CV888-720p to produce the resulting RGB video.
@@
| Output of Cosmos-Predict1-13B-Video2World | |||
| Output of Cosmos-Predict1-13B-Video2World + diffusion decoder | |||